OpenStudy (anonymous):

What is the equation of the line (-4,0) and (0,5) y=????

OpenStudy (campbell_st):

1st step is to find the slope... can you do that..?

OpenStudy (anonymous):

is it m=(y2-y1)/(x2-x1)?

OpenStudy (campbell_st):

thats it so then use y = mx + b to find b, just substitute either point into the equation.

OpenStudy (anonymous):

ive gotten 0=(5/4)x-4+b with b = -5 and 5=(5/4)x(0)+b with b=5

OpenStudy (anonymous):

so is the answer y=(5/4)x+(-1)?

OpenStudy (campbell_st):

well is the slope is 5/4 use x = -4 and y = 0 so 0 = 5/4 * (-4) + b 0 = -5 + b so b = 5 then the equation is y = 5/4x + 5

OpenStudy (anonymous):

OHHH! so i would only need to substitute one equation instead of both?

Directrix (directrix):

You got 5/4 for the slope of the line. The line has both the two given points on it. Choose one of them, say (0,5) Use the point-slope fomat: y - 5 = (5/4) (x - 0) Simplify that.

Directrix (directrix):

(0,5) is the y intercept of the line. You can also use the y = mx+b form. Replace m with the slope 5/4 and b with 5 which is the y-intercept. You'll get the same equation either way. @dada6281

OpenStudy (anonymous):

You are a GENUIS @Directrix